OpenStudy (anonymous):

If the area of a certain triangle is 38 square centimeters and the base is 8 centimeters, what is the height? Please show all of your work.

OpenStudy (saifoo.khan):

Area = 1/2 * base * height. 38 = 0.5 * 8 * h Solve for h.

jimthompson5910 (jim_thompson5910):

Area = (base*height)/2 A = (bh)/2 38 = (8h)/2 38*2 = 8h 76 = 8h 76/8 = h 19/2 = h h = 19/2 h = 9.5

OpenStudy (anonymous):

how do you do this? solve the equation for s v=s+12t over 9

jimthompson5910 (jim_thompson5910):

v=(s+12t)/9 9v=s+12t 9v-12t=s s=9v-12t